______________3. An oxide of chromium is made up of 5.20 g chromium and 5.60 g

oxygen What is the empirical formula of the oxide? (Note: An oxide of nitrogen

contains 63.1% oxygen and has a molar mass of 76.0 g/mol.)

______________4. What is the empirical formula for this compound?

______________5. What is the molecular formula of the compound?


1
Expert's answer
2022-03-23T06:12:01-0400

CrxOy

n(Cr) = m(Cr)/M(Cr) = 5.20 g / 52 g/moll = 0.1 moll

n(O) = m(O)/M(O) = 5.60 g / 16 g/moll = 0.35 moll

n(Cr) : n(O) = 0.1 : 0.35 = 1 : 3.5 = 2 : 7

x=2, y=7


formula of the oxide - Cr2O- Dichromate


molecular formul - Cr2O72-

Dichromate(2-) is a divalent inorganic anion obtained by removal of both protons from dichromic acid (Cr2H2O7)